INAF LBC+MODS queue observing, 2012 January 18-19

Observer: K. Boutsia, R. Speziali
Telescope Operator: S. Allanson
Support Astronomer: O. Kuhn
Instrument: LBC+MODS

Summary

Thin cirrus on the horizon at twilight but most of the sky was clear.
We have not been able to use MODS tonight because of a technical issues . Actually a piece of the inner baffling came undone and was blocking the red channel. ~3hr lost because of this and we had to make significant adjustments to our observing plan.
We switched to LBC but it has been really hard to keep the focus at the beginning because of big temperature changes and the seeing was variable between 1.2''-1.6''.
We did use the focus.obs and we were doing lbcrangebal before each new target.
We had to perform extra collimation in the middle of OBs, and FPIA would take longer than expected in some instances.
When we finally reached thermal equilibrium the images were round and the seeing stabilized <1.2'', with parts of the night <1.0''.

Data Summary

Project Objects Filter ExpTime Notes
Giallongo RXJ0926 g-Sloan / r-Sloan 1hr problems with collimation at the beginning - it should be within specs
Piotto WDs_M67 B-bess / i-Sloan 2hr most of the image have <1.1'' seeing. The project was finished at airm. 1.5 (PI request 1.2)
Sollima Pal4cluster g-Sloan / r-Sloan 1hr most of the exposures have <1.0'' seeing.
PHOT_STD SA98 UsgrBrizy   airmass 1.3''
PHOT_STD P309 UBg/riz   airmass 1.1''

Details

00:52 Starting up MODS
  • mods1 start gui
  • execMODS ~MODSeng/modsScripts/Support/modsWake.pro
  • modsDisp

01:25 The dimm gives 1.5'' seeing ... we have trouble collimating close to mac0647 ( that is relatively low right now, 35elev...), so we go to a standard star that is in higher elevation (the one that Radovich uses)..

01:29 acqMODS g191b2b.acq mods1r 0001-0004
  • dx=1.084, dy=8.119

01:42 execMODS g191b2b.obs mods1b 0001 / mods1r 0005

01:51 acqMODS macsj0647_mods.acq mods1r 0006-0007
  • there is some dark structure obsuring the upper left corner of the image and it is not the guide probe.
  • we do an offset to try and understand what this is... of course then we will need to repeat the acquisition mods1r 0008

02:10 calling Rick -MODS is KO for tonight - we cannot use it!

3:50 switching to LBC

04:12 Slew to ACT0318 for pointing correction and collimation
  • dofpia
  • dimm seeing 1.5''
  • lbcrangbal

04:27 P309U_all - standard star

04:44 repeat P309U_all with all necessary filters (airmass 1.12)

04:53 SA98_UsgrBrizy (aimass 1.3)

05:07 Giallongo_RXJ0926_focus
  • for dofpia and lbcrangebal

05:16 Giallongo_RXJ0926_GR1 (airmass - 1.65... the PI asks for 1.3)
  • seeing on the images: blue 1.6'' / red 1.35''
  • according to the ETC between airmass 1.6 and 1.3 you expect to loose 0.07mag in the gsloan and 0.02mag in the rsloan... so we keep going
  • seeing is marginally within spec for the blue and well within spec for the red
  • seeing on the images: blue 1.4'' / red 1.3''

05:58 dofpia

06:02 Giallongo_RXJ0926_GR2 * seeing on the images: blue 1.2'' / red 1.1''

06:35 abort OB because images are out of focus... there is a lot of coma, especially in the red band.
  • dofpia
  • since the last time we collimated the temp has changed 1.5deg
  • by the time the red was converged the blue developed coma
  • repeat dofpia,/X2

06:45 Giallongo_RXJ0926_GR2
  • seeing on the images is ~1.4''

06:55 abort again because the blue images are elongated
  • dofpia,/X2
  • and call Andrew

7:02 Giallongo_RXJ0926_GR2

7:40 slew to Piotto for collimation and lbcrangebal
  • WDs_M67_focus.ob

08:02 WDs_M67_LONGBi1
  • seeing on images ~1.1'' but dimm 1.4''
  • the blue is shifting out of focus!!!!
  • the seeing on the blue side is 1.4'' and ~1.1 on the red

*08:20 OB stopped because the blue channel is loosing focus. It was 1.4". Red was 1".
  • The blue channel is not sending AO corrections since the stars in the tech chip are too faint.
  • Both tech chips have stars though...
  • dofpia - it took 5!!! iterations to converge!

08:27 WDs_M67_LONGBi1
  • first image both channels ~1.0'' seeing
  • dimm 1.2''
  • now also the blue is sending corrections

08:53 dofpia

08:58 WDs_M67_LONGBi2
  • seeing on the images ~1.0''
  • dimm seeing 1.2''

09:23 dofpia

09:26 WDs_M67_LONGBi3
  • seeing on the image ~1.1''

09:53 dofpia

09:57 WDs_M67_LONGBi4
  • this OB had the red filter black and we changed it putting i-sloan as for the rest

10:00 WDs_M67_LONGBi4
  • with the i-sloan for the red channel
  • seeing in both channels ~1.0''
  • it ends with a seeing of ~1.1 on the images

10:26 dofpia
  • the blue channel was more out of focus than the red...

10:30 WDs_M67_LONGBi5
  • seeing in the image ~1.0''

10:55 dofpia

10:57 WDs_M67_LONGBi6
  • airmass ~1.45 at the end of the OB but seeing <1.0''

11:23 dofpia

11:26 WDs_M67_MIDBi

11:34 WDs_M67_SHORTBi
  • the program finished at 1.55 airmass and the seeing was <0.9'' for these last exposures
  • we hope they will consider it finished!!

11:44 Pal4focus
  • dofpia and lbcrangebal
  • it is not converging - after 7 iterations we exit dofpia and double the exptime
  • dofpia,/X2

12:00 Pal4cluster_a
  • seeing measured on images <0.9''

12:34 dofpia

12:38 Pal4cluster_b
  • seeing getting worse ~1.2'' but we are in twilight

01:15 Slew for flats

01:36 counts start to be high enough for flats
  • g-sloan / i-sloan - PA=0, 1sec/1sec (x10) - counts 10000 to 40000
  • g-sloan / i-sloan - PA=180, 1sec/1sec (x6) - counts - too high only 1 image obtained
  • B-bessel / r-sloan - PA=180, 1sec/1sec (x2) - counts 15000 - 30000
  • B-bessel / r-sloan - PA= 0, 1sec/1sec (x1.5) - counts >30000

01:58 End of night
  • 10Bias_Bino

-- KonstantinaBoutsia - 19 Jan 2012
Topic revision: r4 - 20 Jan 2012, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback